# PULL: PU-Learning-based Accurate Link Prediction

### v1

@@ 3       $\theta\leftarrow\theta^{\mathrm{new}}$;
        // Equations (3), (4)
4       Approximate $\bar{\mathcal{G}}$ to $\bar{\mathcal{G}}^{\prime}$ by keeping edges with high confidence, while removing the rest;
@@
3.2. Expectation of Graph Structure
@@
During the training process of a link predictor $f_{\theta}$, PULL propagates information through the expected graph structure $\bar{\mathcal{G}}$ of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ over the latent variable $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ where $\theta$ is the learnable model parameter. $\bar{\mathcal{G}}$ requires computing the joint probabilities $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ for all possible graph structures $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. This is intractable since there are $2^{\mid\mathcal{E}_{\mathcal{U}}\mid}$ possible states of $\mathbf{z}$ in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. Instead, PULL efficiently computes the expected graph structure by carefully designing the joint probability $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$.
@@
We convert the graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ with latent variables $\mathbf{z}$ into a line graph $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))=(\mathcal{V}_{L},\mathcal{E}_{L})$ where nodes in $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$ represent the edges of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$, and two nodes in $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$ are connected if their corresponding edges in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ are adjacent. Note that $\mathcal{V}_{L}$ contains both $\mathcal{E}_{\mathcal{P}}$ and $\mathcal{E}_{\mathcal{U}}$ of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ since every node pair $(i,j)$ in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ is correlated with variable $z_{ij}$. We then consider the line graph as a pairwise Markov network, which assumes that any two random variables in the network are conditionally independent of each other given the rest of the variables if they are not directly connected (Parsons, 2011). This simplifies the probabilistic modeling on graph-structured random variables, and effectively marginalizes the joint distribution of nodes in $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$, which corresponds to the distribution $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ of edges in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$.
@@
With the Markov property, the joint distribution of nodes in the line graph $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))=(\mathcal{V}_{L},\mathcal{E}_{L})$ is computed by the multiplication of all the node and edge potentials:
(1) $\small p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)=\frac{1}{F}\prod_{ij\in\mathcal{V}_{L}}\phi_{ij}(z_{ij}\mid\mathbf{X},\theta)\prod_{(ij,jk)\in\mathcal{E}_{L}}\psi_{ij,jk}(z_{ij},z_{jk}\mid\mathbf{X},\theta)$
@@
where $\phi_{ij}$ and $\psi_{ij,jk}$ are node and edge potentials for each transformed node $ij$ and edge $(ij,jk)$, respectively. The node potential $\phi_{ij}$ represents the unnormalized marginal linking probability between nodes $i$ and $j$ in the original graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. The edge potential $\psi_{ij,jk}$ denotes a degree of homophily between the edges containing a common node in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. $F$ is the normalizing factor that ensures the distribution adds up to one. For simplicity, we omit $\mathbf{X}$ in $\phi_{ij}$ and $\psi_{ij,jk}$ in the rest of the paper.
@@
where $h_{i}$ is the hidden representation of node $i$ in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ parameterized by $\theta$, and $\phi_{ij}(z_{ij}=0\mid\theta)=1-\phi_{ij}(z_{ij}=1\mid\theta)$. We set $\phi_{ij}(z_{ij}=1\mid\theta)=1$ for $e_{ij}\in\mathcal{E}_{\mathcal{P}}$ since the linking probability of an observed edge of $\mathcal{G}_{\mathcal{P}}$ is 1. We define $\psi_{ij,jk}$ as a constant $c$ to make the joint distribution focus on the marginal linking probabilities. Then the normalizing constant $F$ in Equation (1) becomes $c^{|\mathcal{E}_{L}|}$ since $\sum_{\mathbf{z}}\prod_{ij\in\mathcal{V}_{L}}\phi_{ij}(z_{ij}\mid\theta)=1$ (see Lemma 1 in Appendix A for proof).
@@
As a result, the joint probability $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ is expressed by the multiplication of node potentials of the line graph:
(2) $\begin{split}p&(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)=\prod_{ij\in\mathcal{V}_{L}}\phi_{ij}(z_{ij}\mid\theta)\\ &=\prod_{e_{ij}\in(\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{U}})}\phi_{ij}(z_{ij}\mid\theta)=\prod_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\phi_{ij}(z_{ij}\mid\theta).\end{split}$
@@
Using the marginalized joint probability $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ in Equation (2), we express the expected graph structure $\mathcal{\bar{G}}$ with regard to the latent variables $\mathbf{z}$. Let $\mathbf{A}(\mathbf{z})$ be the adjacency matrix representing the state $\mathbf{z}$ where the $(i,j)$-th component of $\mathbf{A}(\mathbf{z})$, which we denote as $\mathbf{A}(\mathbf{z})_{ij}$, is $z_{ij}\in\{1,0\}$. Then the corresponding weighted adjacency matrix $\mathbf{A}^{\mathcal{\bar{G}}}$ of the expected graph $\mathcal{\bar{G}}$ is computed as follows:
(3) $\begin{split}\mathbf{A}^{\mathcal{\bar{G}}}=\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)}[\mathbf{A}(\mathbf{z})]&=\sum_{\mathbf{z}}p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)\mathbf{A}(\mathbf{z})\\ &=\sum_{\mathbf{z}}\prod_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\phi_{ij}(z_{ij}\mid\theta)\mathbf{A}(\mathbf{z}).\end{split}$
The $(i,j)$-th component $\mathbf{A}^{\mathcal{\bar{G}}}_{ij}$ of $\mathbf{A}^{\mathcal{\bar{G}}}$ is simply expressed as follows:
@@
(4) $\begin{split}\mathbf{A}^{\mathcal{\bar{G}}}_{ij}&=\phi_{ij}(z_{ij}=1\mid\theta)\sum_{\mathbf{z}\mid z_{ij}=1}\prod_{e_{kl}\in\mathcal{E}_{\mathcal{U}}\setminus\{e_{ij}\}}\phi_{kl}(z_{kl}\mid\theta)\mathbf{A}(\mathbf{z})_{ij}\\ &=\phi_{ij}(z_{ij}=1\mid\theta)\end{split}$
@@
since $\mathbf{A}(\mathbf{z})_{ij}=1$ for $z_{ij}=1$, and $\sum_{\mathbf{z}\mid z_{ij}=1}\prod_{e_{kl}\in\mathcal{E}_{\mathcal{U}}\setminus\{e_{ij}\}}\phi_{kl}(z_{kl}\mid\theta)=1$ (see Lemma 1 in Appendix A for proof). As a result, we simply express the expected graph $\mathcal{\bar{G}}$ by an weighted adjacency matrix $\mathbf{A}^{\mathcal{\bar{G}}}$ where $\mathbf{A}^{\mathcal{\bar{G}}}_{ij}=\phi_{ij}(z_{ij}=1\mid\theta)$.
@@
Using the expected graph $\mathcal{\bar{G}}$ directly to train the link predictor $f$ may lead to oversmoothing problem, as $\mathcal{\bar{G}}$ is a fully connected graph represented by $\mathbf{A}^{\mathcal{\bar{G}}}$. Moreover, the training time increases exponentially with the number of nodes. To address these challenges, PULL utilizes an approximated one of $\mathcal{\bar{G}}$ for training $f$, which contains edges with high confidence. Specifically, we approximate $\mathcal{\bar{G}}$ by keeping the top-$K$ edges with the largest weights, while removing the rest. This effectively reduces the number of message passing operation for each layer of $f$ from $\mathcal{V}^{2}$ to $\mathcal{E}+K$ where $\mathcal{V}$ and $\mathcal{E}$ are sets of nodes and edges in the given graph $\mathcal{G}_{\mathcal{P}}$. We denote this approximated one as $\bar{\mathcal{G}}^{\prime}$, and its adjacency matrix as $\mathbf{A}^{\mathcal{\bar{G}}^{\prime}}$.
@@
Another challenge lies in the need to compute weights for every node pairs in each outer iteration of Algorithm 1 to acquire the top-$K$ edges, which results in computational inefficiency. To address this, we define a set of candidate edges determined by the node degrees. This stems from the observation that in real world networks, nodes with higher degrees exhibit a greater likelihood of forming new connections (preferential attachment) (Barabási and Albert, 1999). The candidate edge set consists of node pairs where at least one node has top-$M$ degree among all the nodes. We set $M=100$ in our experiments. PULL selects top-$K$ edges among the candidate edge set instead of all node pairs to approximate $\mathcal{\bar{G}}$.
We gradually increase the number $K$ of selected edges in proportion to that of observed edges through the iterations, which is expressed by $K\leftarrow K+r|\mathcal{E}_{\mathcal{P}}|$, giving more trust in the expected graph structure $\bar{\mathcal{G}}$. This is because the quality of $\bar{\mathcal{G}}$ improves through the iterations (see Figure 2). We set $r=0.05$ in our experiments.
3.3. Iterative Learning
@@
At each iteration, PULL computes the expected graph $\bar{\mathcal{G}}$ given a trained link predictor $f_{\theta}$ with current parameter $\theta$. Then PULL propagates information through $\bar{\mathcal{G}}^{\prime}$, instead of the given edge-incomplete graph $\mathcal{G}_{\mathcal{P}}$ to train a new link predictor $f_{\theta}$ with new parameter $\theta^{\mathrm{new}}$. This prevents PULL from overfitting to $\mathcal{G}_{\mathcal{P}}$, thus improving the link prediction performance.
@@
To optimize the new parameter $\theta^{\mathrm{new}}$, we propose the binary cross entropy loss $\mathcal{L}_{E}$ in Equation (5) by treating the given edges in $\mathcal{E}_{\mathcal{P}}$ and the unconnected edges in $\mathcal{E}_{\mathcal{U}}$ as positive and unlabeled (PU) examples, respectively. For the unconnected edges, we use the expected linking probabilities $\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij}$, which are obtained from the current link predictor $f_{\theta}$, as pseudo labels for $e_{ij}$:
@@
(5) $\begin{split}\mathcal{L}_{E}&=-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\hat{y}_{ij}-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{U}}^{r}}\log(1-\hat{y}_{ij})\\ &\quad-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}^{r}}\bigl{(}\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij}\log\hat{y}_{ij}+(1-\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij})\log(1-\hat{y}_{ij})\bigr{)}\end{split}$
@@
where $\hat{y}_{ij}=f_{\theta^{\mathrm{new}}}(i,j)$. $\mathcal{E}_{\mathcal{P}}^{r}$ is the set of relatively connected edges selected from $\mathcal{E}_{\mathcal{U}}$ when approximating the expected graph structure $\mathcal{\bar{G}}$ by $\mathcal{\bar{G}}^{\prime}$ in Section 3.2, and $\mathcal{E}_{\mathcal{U}}^{r}=\mathcal{E}_{\mathcal{U}}\setminus\mathcal{E}_{\mathcal{P}}^{r}$.
@@
However, in real-world graphs, there is a severe imbalance between the numbers of connected edges and unconnected ones. We balance them by randomly sampling $|\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{P}}^{r}|$ unconnected edges among $\mathcal{E}_{\mathcal{U}}^{r}$ for every epoch. Then Equation (5) is written as follows:
(6) $\begin{split}\mathcal{L}^{\prime}_{E}=&-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\hat{y}_{ij}-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{U}}^{\prime}}\log(1-\hat{y}_{ij})\\ &-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}^{r}}\bigl{(}\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij}\log\hat{y}_{ij}+(1-\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij})\log(1-\hat{y}_{ij})\bigr{)}\end{split}$
where $\mathcal{E}_{\mathcal{U}}^{\prime}$ is the set of randomly sampled edges among $\mathcal{E}_{\mathcal{U}}^{r}$ with the size $|\mathcal{E}_{\mathcal{U}}^{\prime}|=|\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{P}}^{r}|$.
@@
If the current parameter $\theta$ of the link predictor is inaccurate, the quality of the expected graph structure deteriorates, leading to the next iteration’s parameter $\theta^{\mathrm{new}}$ becoming even more inaccurate. Thus, we propose another loss term $\mathcal{L}_{C}$ for correction, which measures the binary cross entropy for all observed edges and randomly sampled unconnected edges from $\mathcal{E}_{\mathcal{U}}^{r}$:
(7) $\mathcal{L}_{C}=-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\tilde{y}_{ij}-\sum_{e_{kl}\in\mathcal{E}_{\mathcal{U}}^{\prime\prime}}\log(1-\tilde{y}_{ij})$
@@
where $\mathcal{E}_{\mathcal{U}}^{\prime\prime}$ is the set of randomly sampled node pairs from $\mathcal{E}_{\mathcal{U}}^{r}$ with size $|\mathcal{E}_{\mathcal{U}}^{\prime\prime}|=|\mathcal{E}_{\mathcal{P}}|$. $\tilde{y}_{ij}$ is $f_{\theta^{\mathrm{new}}}(i,j)$ computed with the given graph $\mathcal{G}_{\mathcal{P}}$ instead of $\bar{\mathcal{G}}$. $\mathcal{L}_{C}$ effectively prevents excessive self-reinforcement in the link predictor of PULL (see Figure 3).
@@
As a result, PULL finds the best parameter $\theta^{\mathrm{new}}$ for each iteration by minimizing the sum of the two loss terms in Equations (6) and (7). We denote the final loss function as $\mathcal{L}(\theta^{\mathrm{new}};\mathcal{\bar{G}}^{\prime},\mathbf{X})=\mathcal{L}_{E}^{\prime}+\mathcal{L}_{C}$. The new parameter $\theta^{\mathrm{new}}$ is used as the current $\theta$ for the next iteration. The iterations stop if the maximum number of iterations is reached or the early stopping condition (see Section 4.1) is met.
3.4. Theoretical Analysis
@@
We theoretically analyze PULL in terms of its connection to the EM (Expectation-Maximization) algorithm, and the time complexity.
Connection to EM algorithm

### v2

@@ 3       $\theta\leftarrow\theta^{\mathrm{new}}$;
        // Equations (3, 4)
4       Approximate $\bar{\mathcal{G}}$ to $\bar{\mathcal{G}}^{\prime}$ by selecting $K$ confident edges from a set of candidate edges;
@@
Expectation of Graph Structure (C2)
@@
During training of a link predictor $f_{\theta}$, PULL propagates information through the expected graph $\bar{\mathcal{G}}$ of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ over the latent variable $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ where $\theta$ is the learnable model parameter. Note that $\bar{\mathcal{G}}$ requires computing the joint probabilities $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ for all possible graph structures $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. This is intractable since there are $2^{\mid\mathcal{E}_{\mathcal{U}}\mid}$ possible states of $\mathbf{z}$ in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. Instead, PULL efficiently computes the expectation of graph structure by carefully designing the joint probability $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$. In the following, we describe how we design the joint probability and compute the expectation of graph.
Designing the joint probability.
@@
We convert the graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ into a line graph $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))=(\mathcal{V}_{L},\mathcal{E}_{L})$ where nodes in $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$ represent the edges of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$, and two nodes in $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$ are connected if their corresponding edges in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ are adjacent. $\mathcal{V}_{L}$ contains both $\mathcal{E}_{\mathcal{P}}$ and $\mathcal{E}_{\mathcal{U}}$ of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ since every node pair $(i,j)$ in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ is correlated with variable $z_{ij}$. We then consider the line graph as a pairwise Markov network, which assumes that any two random variables in the network are conditionally independent of each other given the rest of the variables if they are not directly connected (Koller and Friedman 2009). This simplifies the probabilistic modeling on graph-structured random variables, and effectively marginalizes the joint distribution of nodes in $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$, which corresponds to the distribution $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ of edges in the original graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$.
@@
With the Markov property, the distribution $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ is computed by the multiplication of all the node and edge potentials in the line graph $L(\mathcal{G}_{\mathcal{P}}(\mathbf{z}))$:
$p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)=\\ \frac{1}{F}\prod_{ij\in\mathcal{V}_{L}}\phi_{ij}(z_{ij}\mid\mathbf{X},\theta)\prod_{(ij,jk)\in\mathcal{E}_{L}}\psi_{ij,jk}(z_{ij},z_{jk}\mid\mathbf{X},\theta)$ (1)
@@
where $\phi_{ij}$ is the node potential for a transformed node $ij$, and $\psi_{ij,jk}$ is the edge potential for a transformed edge $(ij,jk)$. The node potential $\phi_{ij}$ represents the unnormalized marginal linking probability between nodes $i$ and $j$ in the original graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. The edge potential $\psi_{ij,jk}$ denotes a degree of homophily between the edges containing a common node in $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$. $F$ is the normalizing factor that ensures the distribution adds up to one. For simplicity, we omit $\mathbf{X}$ in $\phi_{ij}$ and $\psi_{ij,jk}$ in the rest of the paper.
@@
where $\phi_{ij}(z_{ij}=0\mid\theta)=1-\phi_{ij}(z_{ij}=1\mid\theta)$ and $f_{\theta}(i,j)$ is the predicted marginal linking probability between nodes $i$ and $j$. We set $\phi_{ij}(z_{ij}=1\mid\theta)=1$ for $e_{ij}\in\mathcal{E}_{\mathcal{P}}$ since the linking probability of an observed edge is 1. We use a GCN followed by a sigmoid function as the link predictor: $f_{\theta}(i,j)=\sigma(h_{i}\cdot h_{j})$ where $h_{i}$ is the hidden representation of node $i$ computed by the GCN embedding function with graph $\bar{\mathcal{G}}$. In the first iteration, we initialize $\bar{\mathcal{G}}$ with $\mathcal{G}_{\mathcal{P}}$. Other graph-based models can also be used instead of GCN (see Q2 of Experiments section). We define $\psi_{ij,jk}$ as a constant $c$ to make the joint distribution focus on the marginal linking probabilities. Then the normalizing constant $F$ in Equation (1) becomes $c^{|\mathcal{E}_{L}|}$ as $\sum_{\mathbf{z}}\prod_{ij\in\mathcal{V}_{L}}\phi_{ij}(z_{ij}\mid\theta)=1$ (see Lemma 1 in Appendix for proof).
@@
As a result, the probability $p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)$ is expressed by the multiplication of node potentials $\phi_{ij}$ for $e_{ij}\in\mathcal{E}_{\mathcal{U}}$:
$\begin{split}p(\mathbf{z}\mid\mathbf{X},\;&\mathcal{E}_{\mathcal{P}},\theta)=\prod_{ij\in\mathcal{V}_{L}}\phi_{ij}(z_{ij}\mid\theta)\\ &=\prod_{e_{ij}\in(\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{U}})}\phi_{ij}(z_{ij}\mid\theta)=\prod_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\phi_{ij}(z_{ij}\mid\theta).\end{split}$ (2)
Computing the expectation of graph.
@@
Let $\mathbf{A}(\mathbf{z})$ be the adjacency matrix representing the state $\mathbf{z}$ where the $(i,j)$-th component of $\mathbf{A}(\mathbf{z})$, which we denote as $\mathbf{A}(\mathbf{z})_{ij}$, is $z_{ij}\in\{1,0\}$. Then the corresponding weighted adjacency matrix $\mathbf{A}^{\mathcal{\bar{G}}}$ of the expected graph $\mathcal{\bar{G}}$ is computed as follows:
$\begin{split}\mathbf{A}^{\mathcal{\bar{G}}}=\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z}\mid\mathcal{E}_{\mathcal{P}},\theta)}[\mathbf{A}(\mathbf{z})]&=\sum_{\mathbf{z}}p(\mathbf{z}\mid\mathbf{X},\mathcal{E}_{\mathcal{P}},\theta)\mathbf{A}(\mathbf{z})\\ &=\sum_{\mathbf{z}}\prod_{e_{ij}\in\mathcal{E}_{\mathcal{U}}}\phi_{ij}(z_{ij}\mid\theta)\mathbf{A}(\mathbf{z}).\end{split}$ (3)
The $(i,j)$-th component $\mathbf{A}^{\mathcal{\bar{G}}}_{ij}$ of $\mathbf{A}^{\mathcal{\bar{G}}}$ is expressed as
@@
$\begin{split}&\phi_{ij}(z_{ij}=1\mid\theta)\sum_{\scriptscriptstyle\mathbf{z}\mid z_{ij}=1}\prod_{\scriptscriptstyle e_{kl}\in\mathcal{E}_{\mathcal{U}}\setminus\{e_{ij}\}}\phi_{kl}(z_{kl}\mid\theta)\mathbf{A}(\mathbf{z})_{ij}\\ &=\phi_{ij}(z_{ij}=1\mid\theta)\end{split}$ (4)
@@
where the equality holds since $\mathbf{A}(\mathbf{z})_{ij}=1$ for $z_{ij}=1$, and $\sum_{\mathbf{z}\mid z_{ij}=1}\prod_{e_{kl}\in\mathcal{E}_{\mathcal{U}}\setminus\{e_{ij}\}}\phi_{kl}(z_{kl}\mid\theta)=1$ (see Lemma 1 in Appendix for proof). As a result, we simply express the expected graph $\mathcal{\bar{G}}$ by an weighted adjacency matrix $\mathbf{A}^{\mathcal{\bar{G}}}$ where $\mathbf{A}^{\mathcal{\bar{G}}}_{ij}=\phi_{ij}(z_{ij}=1\mid\theta)$.
Approximating the expected graph.
@@
Propagating information through $\mathcal{\bar{G}}$ directly may lead to oversmoothing problem, as $\mathcal{\bar{G}}$ is a fully connected graph represented by $\mathbf{A}^{\mathcal{\bar{G}}}$. Moreover, the training time increases exponentially with the number of nodes. To address these challenges, PULL utilizes an approximated one of $\mathcal{\bar{G}}$, which contains edges with high confidence. Specifically, we keep the top-$K$ edges with the largest weights while removing the rest. We denote this approximated one as $\bar{\mathcal{G}}^{\prime}$, and its adjacency matrix as $\mathbf{A}^{\mathcal{\bar{G}}^{\prime}}$. From the perspective of PU learning, selecting edges in $\mathcal{\bar{G}}$ can be viewed as selecting relatively connected edges among the unlabeled ones, while treating the rest as relatively unconnected edges. We gradually increase the number $K$ of selected edges in proportion to that of observed edges through the outer iteration of Algorithm 1, which is expressed by $K\leftarrow K+r|\mathcal{E}_{\mathcal{P}}|$, giving more trust in the expected graph $\bar{\mathcal{G}}$. This is because the quality of $\bar{\mathcal{G}}$ improves through the iterations (Figure 3). We set $r=0.05$ in our experiments.
@@
Another challenge lies in the need to compute weights for every node pair in each outer iteration to acquire the top-$K$ edges, which results in computational inefficiency. To address this, we define a set of candidate edges determined by the node degrees. This stems from the observation that nodes with higher degrees exhibit a greater likelihood of forming new connections in real-world networks (Barabási and Albert 1999). The candidate edge set consists of node pairs where at least one node has top-$M$ degree among all the nodes. We set $M=100$ in our experiments. PULL selects top-$K$ edges among the candidate edge set instead of all node pairs to approximate $\mathcal{\bar{G}}$.
Iterative Learning of Link Predictor (C3)
@@
At each iteration, PULL constructs the expected graph $\bar{\mathcal{G}}$ given a trained link predictor $f_{\theta}$ with current parameter $\theta$. Then PULL propagates information through $\bar{\mathcal{G}}^{\prime}$, which is the approximated version of $\bar{\mathcal{G}}$, to train a new link predictor $f_{\theta^{\mathrm{new}}}$ with new parameter $\theta^{\mathrm{new}}$. This effectively prevents the link predictor from blindly trusting the given edge-incomplete $\mathcal{G}_{\mathcal{P}}$.
@@
To optimize the new parameter $\theta^{\mathrm{new}}$, we propose the binary cross entropy (BCE) loss $\mathcal{L}_{E}$ in Equation (5) by treating the given edges in $\mathcal{E}_{\mathcal{P}}$ and the unconnected edges in $\mathcal{E}_{\mathcal{U}}$ as positive and unlabeled (PU) examples, respectively. For the unconnected edges, we use the expected linking probabilities $\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij}$ as pseudo labels for $e_{ij}$:
@@
$\begin{split}\mathcal{L}_{E}&=-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\hat{y}_{ij}-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{U}}^{r}}\log(1-\hat{y}_{ij})\\ &\quad-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}^{r}}\bigl{(}\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij}\log\hat{y}_{ij}+(1-\mathbf{A}^{\mathcal{\bar{G}^{\prime}}}_{ij})\log(1-\hat{y}_{ij})\bigr{)}\end{split}$ (5)
@@
where $\hat{y}_{ij}=f_{\theta^{\mathrm{new}}}(i,j)$. $\mathcal{E}_{\mathcal{P}}^{r}$ is the set of relatively connected edges selected from $\mathcal{E}_{\mathcal{U}}$ when approximating the expected graph structure $\mathcal{\bar{G}}$ by $\mathcal{\bar{G}}^{\prime}$, and $\mathcal{E}_{\mathcal{U}}^{r}=\mathcal{E}_{\mathcal{U}}\setminus\mathcal{E}_{\mathcal{P}}^{r}$.
@@
However, in real-world graphs, there is a severe imbalance between the numbers of connected edges and unconnected ones. We balance them by randomly sampling $|\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{P}}^{r}|$ unconnected edges among $\mathcal{E}_{\mathcal{U}}^{r}$ for every epoch. We denote the loss $\mathcal{L}_{E}$ with a set $\mathcal{E}_{\mathcal{U}}^{\prime}$ of randomly sampled edges among $\mathcal{E}_{\mathcal{U}}^{r}$ with size $|\mathcal{E}_{\mathcal{U}}^{\prime}|=|\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{P}}^{r}|$ as $\mathcal{L}^{\prime}_{E}$.
@@
If the current parameter $\theta$ of the link predictor is inaccurate, the quality of the expected graph structure deteriorates, leading to the next iteration’s parameter $\theta^{\mathrm{new}}$ becoming even more inaccurate. Thus, we propose another loss term $\mathcal{L}_{C}$ for correction, which measures the BCE for all observed edges and randomly sampled unconnected edges from $\mathcal{E}_{\mathcal{U}}^{r}$:
$\mathcal{L}_{C}=-\sum_{e_{ij}\in\mathcal{E}_{\mathcal{P}}}\log\tilde{y}_{ij}-\sum_{e_{kl}\in\mathcal{E}_{\mathcal{U}}^{\prime\prime}}\log(1-\tilde{y}_{ij})$ (6)
@@
where $\mathcal{E}_{\mathcal{U}}^{\prime\prime}$ is the set of randomly sampled node pairs from $\mathcal{E}_{\mathcal{U}}^{r}$ with size $|\mathcal{E}_{\mathcal{U}}^{\prime\prime}|=|\mathcal{E}_{\mathcal{P}}|$. $\tilde{y}_{ij}$ is $f_{\theta^{\mathrm{new}}}(i,j)$ computed with the given graph $\mathcal{G}_{\mathcal{P}}$. $\mathcal{L}_{C}$ effectively prevents excessive self-reinforcement in the link predictor of PULL (Figure 3).
@@
As a result, PULL finds the best parameter $\theta^{\mathrm{new}}$ for each iteration by minimizing the sum of the two loss terms $\mathcal{L}^{\prime}_{E}$ and $\mathcal{L}_{C}$. We denote the final loss function as $\mathcal{L}(\theta^{\mathrm{new}};\mathcal{\bar{G}}^{\prime},\mathbf{X})=\mathcal{L}_{E}^{\prime}+\mathcal{L}_{C}$. The new parameter $\theta^{\mathrm{new}}$ is used as the current $\theta$ for the next iteration. The iterations stop if the maximum number of iterations is reached or an early stopping condition is met.
Theoretical Analysis
@@
We theoretically analyze PULL in terms of its connection to the EM algorithm, and the time complexity.
Relation of PULL to EM algorithm.